# Comparing Two Proxy Methods for Causal Identification

### v2

@@ 4.0.1 Discrete Model Comparison
Assumptions 6 and 7 from Miao, Geng and Tchetgen Tchetgen (2018) do not imply that $2|\mathcal{U}|+2\leq k_{A}+k_{B}+k_{C}$, nor does the converse hold. In other words, the models are non-nested.
4.0.2 Continuous Model Comparison
Table 2 reviews the assumptions of Miao, Geng and Tchetgen Tchetgen (2018) alongside those of Hu and Schennach (2008) for continuous models. We conjecture that the two models are non-nested in the model Markov to Figure 3. We give an example satisfying the assumptions of both approaches, which appears in the latter half of Kuroki and Pearl (2014) and is revisited in Miao, Geng and Tchetgen Tchetgen (2018).
Table 2: Comparison for continuous hidden confounding model Markov to Figure 3.
@@ 4.0.2 Continuous Model Comparison
$W,Z,Y,U\mid A=a$ admit bounded, nonzero densities wrt Lebesgue measure on $\mathcal{W}\!\times\!\mathcal{Z}\!\times\!\mathcal{U}$ and a dominating measure $\mu$ on $\mathcal{Y}$; all marginals and conditionals are bounded
@@
$f_{Y\mid U,A}(y\,|\,u_{i},a)\neq f_{Y\mid U,A}(y\,|\,u_{j},a)$ with positive probability under $Y\mid A=a,\qquad i\neq j$
$U$ $=\mu_{U}+\varepsilon_{U},$ $Z$ $=\beta_{0Z}+\alpha_{UZ}U+\varepsilon_{Z},$ $A$ $=\beta_{0A}+\alpha_{UA}U+\alpha_{ZA}Z+\varepsilon_{A},$ $W$ $=\beta_{0W}+\alpha_{UW}U+\varepsilon_{W},$ $Y$ $=\beta_{0Y}+\alpha_{AY}A+\alpha_{UY}U+\varepsilon_{Y},$ with all disturbances mean-zero Gaussian and mutually independent
@@
Bridge equation approach assumptions are from Miao, Geng and Tchetgen Tchetgen (2018).
@@ 5 A Brief Note on Estimation
While our primary focus in this paper is on comparing model assumptions for identification, it is worth noting that Miao, Geng and Tchetgen Tchetgen (2018) provide an illustrative example in which $f_{Y(a)}(y)$ can be derived analytically under location–scale family assumptions, yielding a tractable framework for consistent estimation of causal effects. In their work, the efficient influence function has also been established under further model assumptions, giving rise to regular estimation of causal effects achieving $\sqrt{n}$-consistency and semiparametric efficiency (Cui et al., 2023). In contrast, Hu and Schennach (2008) rely on sieve methods to approximate factors of the full law. Some semiparametric theory has also been developed for estimating causal effects recovered from the array decomposition approach in the context of proxies for hidden treatment, though this literature remains relatively nascent (Zhou and Tchetgen, 2024).
@@ 6 Conclusion and Future Directions
Proxy-based approaches have emerged as a promising solution to identifying causal targets when hidden variables pose an obstacle. Within this framework, two complementary strategies stand out. The bridge equation method identifies causal targets by solving functional equations, while the array decomposition approach leverages unique determination of eigenspaces to uncover latent structures which constitute those targets. Although both methods hinge on assumptions regarding the informativeness of proxy variables for the latent structure, comparing their respective assumptions and areas of overlap offers clearer insight into the contexts in which each method is most appropriately applied.
@@
Appendix A Frontdoor Adjustment Formula
Consider the hidden variable model in Figure 5, where confounders $U$ between the treatment $A$ and outcome $Y$ are unobserved, however a mediator $M$ between $A$ and $Y$ is observed. Here, the counterfactual distribution $f_{Y(a)}(y)$ is identified by the frontdoor formula in Equation 7, provided the following conditions hold for each $a\in\mathcal{A}$ and $m\in\mathcal{M}$:
@@ Appendix A Frontdoor Adjustment Formula
$M(a)\perp\!\!\!\perp A\text{ and }Y(m)\perp\!\!\!\perp M\mid A.$
@@
Appendix B Kuroki and Pearl (2014)
We summarize the identification proof from Kuroki and Pearl (2014), relying on Lemma B.1.
Lemma B.1.
Let $\mathbf{M}\in\mathbb{R}^{n\times n}$ be an invertible square matrix. Suppose $\mathbf{M}$ admits the eigendecomposition
$\mathbf{M}=\mathbf{V}\mathbf{\Lambda}\mathbf{V}^{-1},$
where $\mathbf{V}\in\mathbb{R}^{n\times n}$ is the matrix whose columns are the eigenvectors of $\mathbf{M}$, and $\mathbf{\Lambda}\in\mathbb{R}^{n\times n}$ is the diagonal matrix of the corresponding eigenvalues. Then the eigendecomposition of $\mathbf{M}$ is uniquely determined up to rescaling of the eigenvectors if and only if the eigenvalues of $\mathbf{M}$ (i.e., the diagonal entries of $\mathbf{\Lambda}$) are distinct.
@@ Appendix B Kuroki and Pearl (2014)
$\mathbf{P}_{y,W\mid Z,a}=\mathbf{P}_{W\mid U}\operatorname{diag}(\mathbf{P}_{y\mid U,a})\mathbf{P}_{U\mid Z,a}.$ (9)
@@
Then Lemma B.1 allows us to identify the column vectors of $\mathbf{P}_{W\mid U}$ under Assumption 11. In other words, we recover conditional distributions $f_{W\mid U}(w\mid u_{i})$, up to the labeling of the latent variable $U$. Inverting $f_{Y,W\mid A}(y,w\mid a)=\sum_{i}f_{Y,U\mid A}(y,u_{i}\mid a)\,f_{W\mid U}(w\mid u_{i}),$ we recover
$f_{Y\mid U,A}(y\mid u_{i},a),\quad f_{U}(u_{i}\mid a).$
Further by $W\perp U\mid A$, latent factors for label $i$ recovered across values of $a\in\mathcal{A}$ can be aligned, giving the components comprising the adjustment formula in Equation (1). ∎
Appendix C Mediator Proxy Analogue for Kuroki and Pearl (2018)
@@
Define for each $a\in\mathcal{A}$,
$\mathbf{P}_{Z\mid M,a}=\big[\,f_{Z\mid M,A}(z\mid m_{1},a),\;\dots,\;f_{Z\mid M,A}(z\mid m_{|\mathcal{M}|},a)\,\big],$
$\mathbf{P}_{W\mid M}=\big[\,f_{W\mid M}(w\mid m_{1}),\;\dots,\;f_{W\mid M}(w\mid m_{|\mathcal{M}|})\,\big],$
$\mathbf{P}_{Y\mid M,a}=\big[\,f_{Y\mid M,A}(y\mid m_{1},a),\;\dots,\;f_{Y\mid M,A}(y\mid m_{|\mathcal{M}|},a)\,\big].$
@@
$W,Z,Y,\text{ are mutually independent }\mid M,A.$
@@
$f_{Y(a)}(y)$ is identified by the proof in Appendix B, replacing $U$ with $M$.

### v3

@@ 4.1 Discrete Model Comparison
Assumptions 6 and 7 from Miao et al. (2018) do not imply that $2|\mathcal{U}|+2\leq k_{A}+k_{B}+k_{C}$, nor does the converse hold. In other words, the models are non-nested.
4.2 Continuous Model Comparison
Table 2 reviews the assumptions of Miao et al. (2018) alongside those of Hu and Schennach (2008) for continuous models. We conjecture that the two models are non-nested in the model represented by Figure 3. We give an example satisfying the assumptions of both approaches, which appears in the latter half of Kuroki and Pearl (2014) and is revisited in Miao et al. (2018).
Table 2: Comparison for continuous hidden confounding model in Figure 3.
@@ 4.2 Continuous Model Comparison
$(W,Z,Y,U,A)$ admit a joint density that is bounded and nonzero.
@@
The set $\{y:f_{Y\mid U,A}(y\mid u_{i},a)\neq f_{Y\mid U,A}(y\mid u_{j},a)\}$ has positive probability under the marginal of $Y\mid A=a$
$\hskip 0.0pt\begin{aligned} U&=\mu_{U}+\varepsilon_{U},\\ Z&=\beta_{0Z}+\alpha_{UZ}U+\varepsilon_{Z},\\ A&=\beta_{0A}+\alpha_{UA}U+\alpha_{ZA}Z+\varepsilon_{A},\\ W&=\beta_{0W}+\alpha_{UW}U+\varepsilon_{W},\\ Y&=\beta_{0Y}+\alpha_{AY}A+\alpha_{UY}U+\varepsilon_{Y},\end{aligned}$ with all disturbances mean-zero Gaussian and mutually independent
@@
Bridge equation approach assumptions are from Miao et al. (2018).
@@ 5 A Brief Note on Estimation
While our primary focus in this paper is on comparing model assumptions for identification, it is worth noting that Miao et al. (2018) provide an illustrative example in which $f_{Y(a)}(y)$ can be derived analytically under location–scale family assumptions, yielding a tractable framework for consistent estimation of causal effects. In their work, the efficient influence function has also been established under further model assumptions, giving rise to estimation of causal effects achieving semiparametric efficiency (Cui et al., 2024). In contrast, Hu and Schennach (2008) rely on sieve methods to approximate factors of the full law. Some semiparametric theory has also been developed for estimating causal effects recovered from the array decomposition approach in the context of proxies of hidden treatment, though this literature remains relatively nascent (Zhou and Tchetgen Tchetgen, 2024).
@@ 6 Conclusion and Future Directions
Proxy-based approaches have emerged as a promising solution to identifying causal targets when hidden variables pose an obstacle. Within this framework, two complementary strategies stand out. The bridge equation method identifies causal targets by solving functional equations, while the array decomposition approach leverages unique determination of eigenspaces to uncover latent distributions which constitute those targets. Although both methods hinge on assumptions regarding the informativeness of proxy variables for the latent structure, comparing their respective assumptions and areas of overlap offers clearer insight into the contexts in which each method is most appropriately applied.
@@
Appendix A Definitions
Let $V$ and $T$ be discrete random variables with supports
$\mathcal{V}=\{v_{1},\dots,v_{|\mathcal{V}|}\},\qquad\mathcal{T}=\{t_{1},\dots,t_{|\mathcal{T}|}\}.$
For each $a\in\mathcal{A}$, define matrix $\mathbf{P}_{V\mid T,a}\in\mathbb{R}^{|\mathcal{V}|\times|\mathcal{T}|}$ by
$\big(\mathbf{P}_{V\mid T,a}\big)_{ij}:=\Pr(V=v_{i}\mid T=t_{j},A=a).$ (9)
When conditioning on $A$ is not intended, we write matrix $\mathbf{P}_{V\mid T}$, defined analogously by
$\big(\mathbf{P}_{V\mid T}\big)_{ij}:=\Pr(V=v_{i}\mid T=t_{j}).$
Appendix B Frontdoor Adjustment Formula
Consider the hidden variable model in Figure 4, where variables $U$ creating treatment-outcome confounding are unobserved, however a mediator $M$ between treatment $A$ and outcome $Y$ is observed. Here, the counterfactual distribution $f_{Y(a)}(y)$ is identified by the frontdoor formula,
$f_{Y(a)}(y)=\iint f_{Y\mid A,M}(y\mid a^{\prime},m)\,f_{M\mid A}(m\mid a)\,f_{A}(a^{\prime})\,da^{\prime}\,dm,$ (10)
provided the following conditions hold for each $a\in\mathcal{A}$ and $m\in\mathcal{M}$:
@@ Appendix B Frontdoor Adjustment Formula
$M(a)\perp\!\!\!\perp A\text{ and }Y(m)\perp\!\!\!\perp M\mid A,$
@@
Appendix C Kuroki and Pearl (2014)
We summarize the identification proof from Kuroki and Pearl (2014).
@@ Appendix C Kuroki and Pearl (2014)
$\mathbf{P}_{y,W\mid Z,a}=\mathbf{P}_{W\mid U}\operatorname{diag}(f_{y\mid U,a})\mathbf{P}_{U\mid Z,a},$ (12)
where $\operatorname{diag}(f_{y\mid U,a})$ denotes the diagonal matrix with $(i,i)$-th entry $\Pr(Y=y\mid U=u_{i},A=a)$.
@@
Then eigendecomposition allows us to identify the column vectors of $\mathbf{P}_{W\mid U}$ under Assumption 11. In other words, we recover conditional distributions $f_{W\mid U}(w\mid u_{i})$, up to the labeling of the latent variable $U$. Inverting $f_{Y,W\mid A}(y,w\mid a)=\sum_{i}f_{Y,U\mid A}(y,u_{i}\mid a)\,f_{W\mid U}(w\mid u_{i}),$ we recover
$f_{Y\mid U,A}(y\mid u_{i},a),\quad f_{U\mid A}(u_{i}\mid a).$
Further, the component $f_{W\mid U}(w\mid u_{i})=f_{W\mid U}(w\mid u_{i},a)$ for all values $a\in\mathcal{A}$, so that the labels can be aligned across treatment levels, yielding the components in Equation (1).
Appendix D Mediator Proxy Analogue for Kuroki and Pearl (2014)
@@
$W,Z,Y\text{ are mutually independent conditional on }M,A.$
@@
$f_{Y(a)}(y)$ is identified by the proof in Appendix C, replacing $U$ with $M$.